OpenStudy (anonymous):

Please Help Me!! The equation of a given circle in general form is x2+ y2 − 8x + 12y + 27 = 0. Write the equation in standard form, (x − h)2 + (y - k)2 = r2, by completing the squares in the equation.

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

first step to make it easier is to write \[x^2-8x+y^2+12y=-27\]

OpenStudy (misty1212):

what is half of 8?

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

4

OpenStudy (misty1212):

ok so we go right to \[(x-4)^2+y^2+12y=-24+4^2\] that takes care of the \(x\) term now what is half of 12?

OpenStudy (anonymous):

6

OpenStudy (mathmale):

Cassandra: Just so that you know, the correct approach is "completing the square." You'll need to know this method by name in the future.

OpenStudy (misty1212):

good, so repeating with the y term we get \[(x-4)^2+(y+6)^2=-27+4^2+6^6\] of course you still have to compute the right hand side of the equal sign this is how you complete the square

OpenStudy (anonymous):

okay thank you very much

OpenStudy (misty1212):

\[\color\magenta\heartsuit\]